The school that Beth goes to is selling tickets to a spring musical. On the first day of ticket sales the school sold
7 adult tickets and 4 student tickets for a total of $165. The school took in $105 on the second day by selling 3
adult tickets and 4 student tickets. Find the price of an adult ticket and the price of a student ticket.

Answer by Boreal(15235)

A is cost of adult ticket
S is cost of student ticket
7A+4S=165
3A+4S=105
==========
7A+4S=165
-3A-4S= -105, multiplying by -1
4A=60
A=$15
===
first equation
105+4S=65
4S=60
S=$15
=====
check with second
45+60=105
=======
Both groups have $15 tickets.